# Antagonistic impact of thermal expansion and phonon anharmonicity on the phonon-limited resistivity of elemental metals from first principles

### v1

@@ Abstract
To include TE effects, we calculate the Helmholtz free energy $F(V,T)$ with both electron and phonon contributions. The electron part is composed of the clamped-ion energy at 0 K and the energy of the thermal electronic excitation. For the phonon contribution, we use the quasi-harmonic approximation (QHA) [40, 41, 42]. We use the same computational parameters as Ref. [27]. We compute $F(V,T)$ from density functional theory (DFT) with Quantum ESPRESSO [43, 44] using a third-order Birch-Murnaghan isothermal fit [45, 46], which produces the temperature-dependent lattice parameter. To include anharmonic effects, we use TDEP [34] where the Hamiltonian is expanded to the second-order with respect to the atom displacements [33]. We use stochastic sampling [47] in which the configurations with atomic displacements are generated by the canonical ensemble from the interatomic force constants (IFCs) and the effective IFCs are determined by linear least-squares solution with the displacements and forces. The IFCs and atomic displacements can therefore be iteratively calculated at different temperatures. Additional anharmonic computational details and convergence studies for Pb are provided in Sec. S1 of the Supplemental Information (SI) [48].
@@
where $\varepsilon_{n\mathbf{k}}$ are the eigenenergies, $f_{n\mathbf{k}}^{0}$ is the equilibrium Fermi-Dirac distribution for electrons, $n_{\mathbf{q}\nu}$ is the Bose-Einstein distribution for phonons, and $\omega_{\mathbf{q}\nu}$ is the frequency of the phonon with the wavevector $\mathbf{q}$ and the branch index $\nu$. The $g_{mn\nu}(\mathbf{k,q})$ are the electron-phonon coupling (EPC) matrix element, which represents the probability amplitudes of the scattering of electron from state $n\mathbf{k}$ to $m\mathbf{k+q}$ by a phonon $\mathbf{q}\nu$. In Eq. (2), we have introduced the electron-phonon scattering rate as [50]
@@
Figure 2: Phonon dispersion of Pb including (a) lattice thermal expansion (TE) effect where calculations are harmonic with a temperature-dependent lattice parameters, (b) without TE and with temperature-dependent phonon anharmonicity, and (c) with TE and phonon anharmonicity. In all cases the gray line is the harmonic phonon dispersion computed with density functional perturbation theory (DFPT) without TE. (d) Phonon dispersion of Pb calculated with TE and phonon anharmonicity at 110 K compared with experimental data at 100 K measured by neutron scattering from Ref. 54 ($\times$ markers) and from Ref. 55 (open and closed circles).
@@
Figure 4: Electrical resistivity of Al from 0 to 1000 K and Nb from 0 to 3000 K using the iterative Boltzmann transport equation (IBTE). The green and blue markers are computed without and with thermal expansion (TE), without considering the phonon anharmonicity, while the computation of orange markers takes the phonon anharmonicity into account. The experimental data are from Ref. [53].
To assess the generality of this antagonistic effect between lattice TE and phonon anharmonicity, we further investigated the electrical resistivity of aluminum and niobium. Unlike Pb, Nb shows a Kohn anomaly related to its nesting Fermi surface [56, 57]. For Al and Nb we use the experimental temperature-dependent lattice constant [58, 59, 60] and provide all computational details and convergence studies in Sec. S3 of the SI [48]. The calculated electrical resistivity matches quite well with experiments below 500 K in Al and Nb when both TE and anharmonicity are excluded, with the comparison shown in Fig. 4. We find that the case of Al is very similar to Pb and that the impact of TE and phonon anharmonicity is opposite and of similar magnitude.
Interestingly, we find that the case of Nb is different, and although the impact of TE and phonon anharmonicity is also opposite, they do not cancel each other. In this case, the effect of the phonon anharmonicity is stronger, which brings the results closer to experiment [53]. We attribute the remaining overestimation to saturation effects [61].

### v2

# Opposite impact of thermal expansion and phonon anharmonicity on the phonon-limited resistivity of elemental metals from first principles

@@ Abstract
To include TE effects, we calculate the Helmholtz free energy $F(V,T)$ with both electron and phonon contributions. The electron part is composed of the clamped-ion energy at 0 K and the energy of the thermal electronic excitation. For the phonon contribution, we use the quasi-harmonic approximation (QHA) [37, 10, 2]. We use the same computational parameters as Ref. [27]. We compute $F(V,T)$ from density functional theory (DFT) with Quantum ESPRESSO [19, 20] using a third-order Birch-Murnaghan isothermal fit [7, 45], which produces the temperature-dependent lattice parameter. To include anharmonic effects, we use TDEP [36] where the Hamiltonian is expanded to the second-order with respect to the atom displacements [32]. We use stochastic sampling [58] in which the configurations with atomic displacements are generated by the canonical ensemble from the interatomic force constants (IFCs) and the effective IFCs are determined by linear least-squares solution with the displacements and forces. The IFCs and atomic displacements can therefore be iteratively calculated at different temperatures. Additional anharmonic computational details and convergence studies for Pb are provided in Sec. S1 of the Supplemental Information (SI) [68] (see also references  [46, 44, 47, 4, 5, 25, 65, 48, 69, 42, 31, 13, 58, 19, 64, 40, 20, 36, 27] therein).
@@
where $\varepsilon_{n\mathbf{k}}$ are the eigenenergies, $f_{n\mathbf{k}}^{0}$ is the equilibrium Fermi-Dirac distribution for electrons, $n_{\mathbf{q}\nu}$ is the Bose-Einstein distribution for phonons, and $\omega_{\mathbf{q}\nu}$ is the frequency of the phonon with the wavevector $\mathbf{q}$ and the branch index $\nu$. The $g_{mn\nu}(\mathbf{k,q})$ are the electron-phonon coupling (EPC) matrix element, which represents the probability amplitudes of the scattering of electron from state $n\mathbf{k}$ to $m\mathbf{k+q}$ by a phonon $\mathbf{q}\nu$. We neglect the $g_{mn\nu}(\mathbf{k,q})$ for which $\omega_{\mathbf{q}\nu}<5$ cm-1. In Eq. (2), we have introduced the electron-phonon scattering rate as [52]
@@
Figure 2: Phonon dispersion of Pb including (a) lattice thermal expansion (TE) effect where calculations are harmonic with a temperature-dependent lattice parameters, (b) without TE and with temperature-dependent phonon anharmonicity, and (c) with TE and phonon anharmonicity. In all cases the gray line is the harmonic phonon dispersion computed with density functional perturbation theory (DFPT) without TE. (d) Phonon dispersion of Pb calculated with TE and phonon anharmonicity at 110 K compared with experimental data at 100 K measured by neutron scattering from Ref. 8 ($\times$ markers) and from Ref. 9 (open and closed circles).
@@
Figure 4: Electrical resistivity of (a) Al from 0 to 1000 K and (b) Nb from 0 to 3000 K using the iterative Boltzmann transport equation (IBTE). The green and blue markers are computed without and with thermal expansion (TE), without considering the phonon anharmonicity, while the computation of orange markers takes the phonon anharmonicity into account. The experimental data are from Ref. [33] and the red line is obtained with the lowest order variational approximation (LOVA) from Ref. [22].
To assess the generality of this antagonistic effect between lattice TE and phonon anharmonicity, we further investigated the electrical resistivity of aluminum and niobium. Unlike Pb, Nb shows a Kohn anomaly related to its nesting Fermi surface [53, 62]. For Al and Nb we use the experimental temperature-dependent lattice constant [4, 46, 69] and provide all computational details and convergence studies in Sec. S3 of the SI [68]. The calculated electrical resistivity matches quite well with experiments below 500 K in Al and Nb when both TE and anharmonicity are excluded, with the comparison shown in Fig. 4. We find that the case of Al is very similar to Pb and that the impact of TE and phonon anharmonicity is opposite and of similar magnitude. We note that there is an underestimation between calculations and experiments above 500 K. A similar underestimation was found in a previous theoretical study using the lowest order variational approximation (LOVA) [22], an approximated solution to the BTE.
Interestingly, we find that the case of Nb is different, and although the impact of TE and phonon anharmonicity is also opposite, they do not cancel each other. In this case, the effect of the phonon anharmonicity is stronger, bringing the results closer to experiment [33]. We attribute the remaining overestimation to saturation effects [30]. The lack of cancellation effects in Nb as compared to Pb and Al can be attributed to the highly dispersive nature of phonon shifts with temperature. When the shifts due to thermal expansion or anharmonicity concentrate in different parts of the Brillouin zone, one can not expect a direct cancellation of the effects on resistivity. We can relate this to the distinctly different origin of the shifts caused by volume change and temperature change. A volume change corresponds to a rigid stretching of bonds whereas at increased temperature Nb sees a smearing of its Fermi surface that screens the nesting vectors. At low temperature, flat sheets of the Fermi surface can be connected with a phonon wavevector which can be identified as the Kohn anomaly along $\Gamma-H$. At elevated temperature, the smearing of the Fermi surface removes the singular behavior and the phonon dispersions straighten [62]. We speculate that this is a general behaviour in metals that have intricate Fermi surface geometry and/or bands in close proximity where temperature effects can bring them into contact with the Fermi surface and alter its topology [71].